# Automatic Network Planning with Digital Radio Twin

### v2

@@ I Introduction
In this paper, we propose AutoPlan, a new automatic network planing framework by leveraging digital radio twin techniques. On the one hand, we derive the DRT with the following steps. First, we crowdsource abundant real-world data from mobile users under existing deployed base stations (e.g., prior deployments or those of other network operators). Second, with the online data, we continuously calibrate the parameters (particularly, the material of buildings) of the existing simulator (i.e., Sionna RT), to reduce the sim-to-real discrepancy. Third, we accelerate the query of the calibrated simulator with GPU parallel computing to improve its tractability. On the other hand, we design a network planning algorithm to derive the optimal base station parameters (particularly, location and power). Specifically, we leverage Bayesian optimization to tackle the non-negligible querying costs of the DRT by creating a Gaussian process as the surrogate model and adopting expected improvement (EI) as the acquisition function. Using the field measurement from Husker-Net (a private 5G network at the University of Nebraska-Lincoln), we extensively evaluate AutoPlan under different deployment scenarios, in terms of both coverage, capacity, and efficiency.
@@
  • The DRT with the attribute of fidelity, tractability, and synchronicity, based on the simulator of Sionna RayTracing;
  • An automatic network planning algorithm, based on Bayesian optimization, to derive optimal base station parameters, assisted by the derived DRT;
@@ II-A Problem
The goal is to derive the DRT to achieve all the attributes of fidelity, tractability, and synchronicity. We consider the scene is given with 3D terrain, including the full set of building objects within the region of interest. Each building is associated with an initial material label, such as marble or concrete, which corresponds to a specific set of material parameters. These material labels serve as the initial configuration for the DRT simulation. Since these default values may not fully reflect the actual propagation environment, a calibration process is subsequently performed to refine the material parameters using real-world signal measurements, thereby improving the fidelity and reliability of the DRT.
In this work, we focus on calibrating two key electromagnetic parameters used in DRT modeling for objects within the region $\mathcal{A}$: conductivity (in S/m) and relative permittivity. Specifically, conductivity characterizes how a material attenuates or reflects incident electromagnetic energy, while relative permittivity affects the bending and delay of waves as they traverse the medium, both playing critical roles in DRT modeling. We use $\boldsymbol{\sigma}=[\sigma_{1},\sigma_{2},\cdots,\sigma_{K}]$ and $\boldsymbol{\epsilon}=[\epsilon_{1},\epsilon_{2},\cdots,\epsilon_{K}]$ to denote the conductivity and relative permittivity of $K$ objects, respectively. The parameter set is thus defined as:
@@ II-B Challenge
Although DRT systems can simulate real-world wireless propagation, discrepancies often arise between their simulated outputs and actual measurements—referred to as the sim-to-real gap. This gap does not stem from a single source, but rather results from the complex interplay of multiple factors, such as multipath effects, material properties, and environmental geometry. Due to this multifactor coupling, the gap cannot be eliminated by simply adjusting a single parameter. The key challenge is to develop a systematic methodology for characterizing and mitigating this complex, coupled discrepancy.
@@ II-C Solution
To address the above problem, we develop a materiel learning method by leveraging the differentiable property of state-of-the-art simulators.
@@
$\mathcal{L}(\boldsymbol{\Theta})=\frac{1}{P}\sum_{p=1}^{P}\left\|\tilde{\mathbf{r}}(\boldsymbol{\Theta})-\hat{\mathbf{r}}\right\|$. (3)
We optimize this loss over $E$ training epochs using stochastic gradient descent [7]. At each epoch $e\in{1,\dots,E}$, the parameters are updated as:
@@
where $\eta$ is the learning rate and $\boldsymbol{\Theta}^{(e)}$ denotes the parameters after the $e$-th epoch. The final optimized parameters $\boldsymbol{\Theta}^{*}$ are obtained after $E$ epochs of training.
@@ III-A Problem
The goal of network planning is to obtain the optimal base station parameters (e.g., location and orientation) that maximizes the specified network performance (e.g., coverage and capacity) in the aforementioned DRT, under a given geographic area.
As illustrated in Fig. 1, we consider a geographic region $\mathcal{A}\subset\mathbb{R}^{2}$ with given 3D terrain, such as buildings, trees, and walkways. Here, we denote $\hat{\mathcal{B}}={(\hat{x}_{m},\hat{y}_{m})}_{m=1}^{M}$ as the deployed $M$ base stations (BSs). To further improve coverage and expand capacity, we aim to deploy an additional $N$ BSs at locations $\bar{\mathcal{B}}={(\bar{x}_{n},\bar{y}_{n})}_{n=1}^{N}$, selected from a feasible region $\mathcal{F}\subseteq\mathcal{A}$ that meets practical placement constraints (e.g., rooftop availability, and height limits). The complete BS set after deployment is $\mathcal{B}=\hat{\mathcal{B}}\cup\bar{\mathcal{B}}$, and the goal is to determine $\bar{\mathcal{B}}$ to maximize overall network performance in $\mathcal{A}$.
@@
Next, we introduce two network performance metrics, including coverage and capacity. We discretize $\mathcal{A}$ into a uniform square grid with resolution $a\times a$ meters, yielding $L=\left\lceil|\mathcal{A}|/{a^{2}}\right\rceil$ sampling points, where $|\mathcal{A}|$ is the area of $\mathcal{A}$, and $\lceil\cdot\rceil$ represents the ceiling function (i.e., rounding up to the nearest integer). For each sampling point $\ell\in{1,\dots,L}$, the received signal characteristics are obtained by using the aforementioned DRT that reflects the real-world propagation environment. The querying of the DRT is governed by a set of tunable parameters $\boldsymbol{\Theta}$, such as material conductivity and permittivity, which influence signal behavior and modelling accuracy.
We evaluate the network coverage by using $r_{\ell}(\boldsymbol{\Theta})$, which denotes the reference signal received power (RSRP) at sampling point $\ell$, corresponding to the strongest signal received from all BSs in the set $\mathcal{B}$. Moreover, we evaluate the network capacity by using the signal-to-noise ratio (SNR), which captures signal quality in the presence of interference. Let $s_{\ell}(\boldsymbol{\Theta})$ represent the SNR at point $\ell$, computed under the current parameter setting $\boldsymbol{\Theta}$ in the DRT. Given the parameter of the DRT, we can obtain RSRP and SNR values at all sampling points, and then define the following performance metrics. Given a predefined RSRP threshold $r_{\mathrm{th}}$, the coverage rate is defined as:
$C(\mathcal{B})=\frac{1}{L}\sum_{\ell=1}^{L}\mathds{1}\left[r_{\ell}(\boldsymbol{\Theta})>r_{\mathrm{th}}\right],$ (5)
@@
$S(\mathcal{B})=\frac{1}{L}\sum_{\ell=1}^{L}\log_{2}\left(1+s_{\ell}(\boldsymbol{\Theta})\right).$ (6)
@@
$T(\mathcal{B})=\alpha\cdot C(\mathcal{B})+S(\mathcal{B}),$ (7)
where $\alpha>0$ is a tunable weight controlling the trade-off between coverage and capacity. Here, a larger $\alpha$ places greater emphasis on coverage, whereas a smaller value prioritizes capacity.
@@ III-C Solution
To address the above problem, we design an automatic network planning method by leveraging Bayesian optimization (BO). BO [8, 9] is the state-of-the-art global searching approach (especially for handling expensive querying costs), generally consisting of a surrogate model and acquisition function. The surrogate model aims to online learn the blackbox function (e.g., mean and variance), where candidates include Gaussian process [10] and Bayesian neural networks [11]. Based on the updated surrogate model, the acquisition function evaluates the customized utility of different actions (i.e., BS locations) in the action space (i.e., $\mathcal{F}$). In each iteration, the next action is selected by maximizing the given acquisition function (e.g., EI and PI) [12].
To support multiple base station deployment, a naive approach in BO is to use the surrogate model to learn the combinational action space (i.e., all candidate BS locations) and overall network performance. However, this naive approach exponentially expands the action space, and practically leads to much slow convergence speed and more cumulative querying costs, in terms of time consumption and computation complexity. Hence, we converts the problem to be a series of incremental base station deployment problems. In other words, we only apply the Bayesian optimization to search for one additional base station deployment at a time. Once the base station location is determined, we put it into the existing deployed base station sets (i.e., $\hat{\mathcal{B}}$), and then create another Bayesian optimization searching for the next base station. This incremental base station stops until the total number of base stations is all deployed.
@@
The next action to query is generally determined by maximizing the selected acquisition function, which defines the utility of all the actions in the action space while balancing exploration and exploitation. Specifically, we adopt the EI as the acquisition function, which quantifies the expected gain over the current best performance $T_{\text{best}}$:
@@
This criterion favours actions that are likely to improve the objective while still exploring regions of high uncertainty. At each iteration $t$, we select the next set of deployment locations by maximizing the acquisition function over the feasible region $\mathcal{F}$:
@@ IV The AutoPlan Algorithm
4Stage 1: Digital radio twin calrbating;
@@ 9   Update parameters $\boldsymbol{\Theta}$ according to (4);
10Set $\boldsymbol{\Theta}^{*}\leftarrow\boldsymbol{\Theta}$;
11
@@ 21   Add $(\bar{b}^{*},T)$ to $\mathcal{D}$;
In this section, we summarize the workflow of the AutoPlan algorithm in Alg. 1. Overall, the algorithm takes the input of 3D terrain of the geographic region and the location of deployed base stations $\hat{\mathcal{B}}$, and generates the output of the location of the $N$ additional base stations. In the stage of digital radio twinning, it observes the real-world RSRPs $\hat{\mathbf{r}}$ from crowdsourced mobile users, and updates the material of buildings. Specifically, it iteratively calculates the loss function by comparing the real-world RSRPs with the simulated RSRPs from the simulator, and performs multiple gradient descent steps. In the stage of network planning, it iteratively search the location of the next base station by 1) selecting a base station location (i.e., maximizing the EI acquisition function); 2) querying the DRT to obtain the network performance; and 3) updating the GP surrogate model with new observation.
@@ V Performance Evaluation
Parameters. We select a $1210~\text{m}\times 1138~\text{m}$ campus area at the University of Nebraska–Lincoln, i.e., the deployment region $\mathcal{A}$. There are $K=127$ objects (i.e., buildings) in this region. We obtain the scene geometry from OpenStreetMap3D and derive the feasible deployment region $\mathcal{F}\subseteq\mathcal{A}$ by extracting the union of building footprints. One existing private 5G base station is located near the southwest corner of $\mathcal{A}$, yielding the initial BS set $\hat{\mathcal{B}}={(\hat{x}_{m},\hat{y}_{m})}_{m=1}^{M}$ with $M=1$. In the region $\mathcal{A}$, we collect $P=1494$ RSRP measurements $\hat{\mathbf{r}}=[\hat{r}_{1},\hat{r}_{2},\ldots,\hat{r}_{P}]$ via drive testing, with a 5G dongle (Quectel RM520N-GL). The accurate location of the reciver is obtained by using RTK-GNSS, providing horizontal accuracy better than $2$ cm. For the deploy base stations, they are with a planar antenna array configured as a $1\times 2$ horizontal linear array, with half-wavelength spacing between elements. The array adopts the 3GPP TR 38.901 directional antenna pattern and vertical polarization, representing a typical 5G small cell configuration. In the evaluation, we aim to select $N=5$ additional BS locations from the feasible region, i.e., $\bar{\mathcal{B}}={(\bar{x}_{n},\bar{y}_{n})}_{n=1}^{5}\subseteq\mathcal{F}$. To calculate coverage, we set $r_{th}=-90\mathrm{dBm}$ and $a=0.2,\mathrm{m}$. The default maximum transmission power of the base station is $43\,\mathrm{dBm}$. Additionally, we set $\alpha=10$ to balance the contributions of coverage and capacity in the target objective. All simulations were conducted on a desktop with an Intel Core i7 CPU, an NVIDIA RTX 3080 GPU, and Ubuntu 24.04 OS, using Sionna RT v1.1.0 [4]. Specifically, we calculate the RSRP and SNR by using RadioMapSolver on a uniform $0.2,\mathrm{m}\times 0.2,\mathrm{m}$ grid. To learn the building materials, we use the Adam optimizer [13] with learning rate $\eta=0.01$, for $E=300$ epochs.
@@ V-A Digital Radio Twinning
In this subsection, we evaluate the DRT for all the attributes of fidelity, tractability, and synchronicity, under the AutoPlan algorithm. Specifically, we collect the field measurement via a drive test in the City Campus of the University of Nebraska-Lincoln, as shown in Fig. 5.
@@
As shown in Fig. 3, the training MSE loss steadily decreases over the 300 gradient descent steps and converges after around 200 steps, with a final minimum value of $113.78$. In other words, the sim-to-real discrepancy is reduced by $11.23(8.98\%)$, by calibrating the parameters of building materials.
@@
In addition, Table I reports the inference time of the DRT under different numbers of deployed base stations. Notably, the average inference time is $0.39\mathrm{s}$ with a single base station and increases to $1.24\mathrm{s}$ when five base stations are deployed, highlighting the computational cost associated with large-scale evaluations.

### v3

@@ I Introduction
In this paper, we propose AutoPlan, a new automatic network planning framework by leveraging digital radio twin techniques. On the one hand, we derive the DRT with the following steps. First, we crowdsource abundant real-world data from mobile users under existing deployed base stations (e.g., prior deployments or those of other network operators). Second, with the online data, we continuously calibrate the parameters (particularly, the material of buildings) of the existing simulator (i.e., Sionna RT), to reduce the sim-to-real discrepancy. Third, we accelerate the query of the calibrated simulator with GPU parallel computing to improve its tractability. On the other hand, we design a network planning algorithm to derive the optimal base station parameters (particularly, location and power). Specifically, we leverage Bayesian optimization to tackle the non-negligible querying costs of the DRT by creating a Gaussian process as the surrogate model and adopting expected improvement (EI) as the acquisition function. Using the field measurement from Husker-Net (a private 5G network at the University of Nebraska-Lincoln), we extensively evaluate AutoPlan under different deployment scenarios, in terms of both coverage, capacity, and efficiency.
@@
  • A digital radio twin (DRT) with the attribute of fidelity, tractability, and synchronicity;
  • An automatic network planning algorithm based on Bayesian optimization, that derives optimal base station parameters assisted by the derived DRT;
@@ II-A Problem
The goal is to derive the DRT to achieve all the attributes of fidelity, tractability, and synchronicity. We consider the scene given with 3D terrain, including the full set of building objects within the region of interest. Each building is associated with an initial material label, such as marble or concrete, which corresponds to a specific set of material parameters. These material labels serve as the initial configuration for the DRT simulation. Since these default values may not fully reflect the actual propagation environment, a calibration process is subsequently performed to refine the material parameters using real-world signal measurements, thus improving the fidelity and reliability of the DRT.
In this work, we focus on calibrating two key electromagnetic parameters used in DRT modeling for objects within the region $\mathcal{A}$: conductivity (in S/m) and relative permittivity. Specifically, conductivity characterizes how a material attenuates or reflects incident electromagnetic energy. At the same time, relative permittivity affects the bending and delay of waves as they traverse the medium, both playing critical roles in DRT modeling. We use $\boldsymbol{\sigma}=[\sigma_{1},\sigma_{2},\cdots,\sigma_{K}]$ and $\boldsymbol{\epsilon}=[\epsilon_{1},\epsilon_{2},\cdots,\epsilon_{K}]$ to denote the conductivity and relative permittivity of $K$ objects, respectively. The parameter set is thus defined as:
@@ II-B Challenge
Although DRT systems can simulate real-world wireless propagation, discrepancies often arise between their simulated outputs and actual measurements—referred to as the sim-to-real gap. This gap does not stem from a single source but rather results from the complex interplay of multiple factors, such as multipath effects, material properties, and environmental geometry. Due to this multifactor coupling, the gap cannot be eliminated by simply adjusting a single parameter. The key challenge is developing a systematic methodology for characterizing and mitigating this complex, coupled discrepancy.
@@ II-C Solution
To address the above problem, we develop a material learning method by leveraging the differentiable property of state-of-the-art simulators.
@@
$\mathcal{L}(\boldsymbol{\Theta})=\frac{1}{P}\sum\nolimits_{p=1}^{P}\left\|\tilde{\mathbf{r}}(\boldsymbol{\Theta})-\hat{\mathbf{r}}\right\|$. (3)
We optimize this loss over $E$ training epochs using stochastic gradient descent [7]. At each epoch $e\in\{1,\dots,E\}$, the parameters are updated as:
@@
where $\eta$ is the learning rate and $\boldsymbol{\Theta}^{(e)}$ denotes the parameters after the $e$-th epoch. In practice, after $E$ epochs of SGD, we use the checkpoint that yields the lowest training loss as our estimate of the optimizer in (2), denoted by $\boldsymbol{\Theta}^{*}$.
@@ III-A Problem
The goal of network planning is to obtain the optimal base station parameters (e.g., location and orientation) that maximize the specified network performance (e.g., coverage and capacity) in the aforementioned DRT, under a given geographic area.
As illustrated in Fig. 1, we consider a geographic region $\mathcal{A}\subset\mathbb{R}^{2}$ with given 3D terrain, such as buildings, trees, and walkways. Here, we denote $\hat{\mathcal{B}}={(\hat{x}_{m},\hat{y}_{m})}_{m=1}^{M}$ as the deployed $M$ base stations (BSs). To further improve coverage and expand capacity, we aim to deploy additional $N$ BSs at locations $\bar{\mathcal{B}}={(\bar{x}_{n},\bar{y}_{n})}_{n=1}^{N}$, selected from a feasible region $\mathcal{F}\subseteq\mathcal{A}$ that meets practical placement constraints (e.g., rooftop availability and height limits). The complete BS set after deployment is $\mathcal{B}=\hat{\mathcal{B}}\cup\bar{\mathcal{B}}$, and the goal is to determine $\bar{\mathcal{B}}$ to maximize the overall performance of the network in $\mathcal{A}$.
@@
Next, we introduce two network performance metrics, including coverage and capacity. We discretize $\mathcal{A}$ into a uniform square grid with resolution $a\times a$ meters, yielding $L=\left\lceil|\mathcal{A}|/{a^{2}}\right\rceil$ sampling points, where $|\mathcal{A}|$ is the area of $\mathcal{A}$, and $\lceil\cdot\rceil$ represents the ceiling function (i.e., rounding up to the nearest integer). For each sampling point $\ell\in\{1,\dots,L\}$, the characteristics of the received signal are obtained using the aforementioned DRT that reflects the real-world propagation environment.
We evaluate network coverage using $r_{\ell}(\boldsymbol{\Theta})$, which denotes the received power of the reference signal (RSRP) at the sampling point $\ell$, corresponding to the strongest signal received from all BSs in the set $\mathcal{B}$. In addition, we evaluate the network capacity using the signal-to-noise ratio (SNR), which captures signal quality in the presence of interference. Let $s_{\ell}(\boldsymbol{\Theta})$ represent the SNR at point $\ell$, computed under the current parameter setting $\boldsymbol{\Theta}$ in the DRT. Given the parameter of the DRT, we can obtain RSRP and SNR values at all sampling points, and then define the following performance metrics. Given a predefined RSRP threshold $r_{\mathrm{th}}$, the coverage rate is defined as:
$C(\mathcal{B})=\frac{1}{L}\sum\nolimits_{\ell=1}^{L}\mathds{1}\left[r_{\ell}(\boldsymbol{\Theta})>r_{\mathrm{th}}\right],$ (5)
@@
$S(\mathcal{B})=\frac{1}{L}\sum\nolimits_{\ell=1}^{L}\log_{2}\left(1+s_{\ell}(\boldsymbol{\Theta})\right).$ (6)
@@
$T(\mathcal{B})=\beta\cdot C(\mathcal{B})+S(\mathcal{B}),$ (7)
where $\beta>0$ is a tunable weight controlling the trade-off between coverage and capacity. Here, a larger $\beta$ places greater emphasis on coverage, whereas a smaller one prioritizes capacity.
@@ III-C Solution
To address the above problem, we design an automatic network planning method by leveraging Bayesian optimization (BO). BO [8, 9] is the state-of-the-art global searching approach (especially for handling expensive querying costs), generally consisting of a surrogate model and acquisition function. The surrogate model aims to online learn the black box function (e.g., mean and variance), where candidates include Gaussian process [10] and Bayesian neural networks [11]. Based on the updated surrogate model, the acquisition function evaluates the customized utility of different actions (i.e., BS locations) in the action space (i.e., $\mathcal{F}$). In each iteration, the next action is selected by maximizing the acquisition function (e.g., expected improvement (EI) or probability of improvement (PI)) [12].
To support multiple base station deployment, a naive approach in BO is to use the surrogate model to learn the combinational action space (i.e., all candidate BS locations) and overall network performance. However, this naive approach exponentially expands the action space and practically leads to a much slower convergence speed and more cumulative querying costs, in terms of time consumption and computation complexity. Hence, we convert the problem to a series of incremental base station deployment problems. In other words, we apply Bayesian optimization only to search for one additional base station deployment at a time. Once the base station location is determined, we put it into the existing deployed base station sets (i.e., $\hat{\mathcal{B}}$), and then create another Bayesian optimization searching for the next base station. This incremental deployment continues until all $N$ base stations are deployed.
@@
The next action to query is generally determined by maximizing the selected acquisition function, which defines the utility of all actions in the action space while balancing exploration and exploitation. Specifically, we adopt the EI as the acquisition function, which quantifies the expected gain over the current best performance $T_{\text{best}}$:
@@
This criterion favors actions that are likely to improve the objective while still exploring regions of high uncertainty. At each iteration $t$, we select the next set of deployment locations by maximizing the acquisition function over the feasible region $\mathcal{F}$:
@@ IV The AutoPlan Algorithm
In this section, we summarize the workflow of the AutoPlan algorithm in Alg. 1. Overall, the algorithm takes the input of 3D terrain of the geographic region and the location of deployed base stations $\hat{\mathcal{B}}$, and generates the output of the location of the $N$ additional base stations. In the digital radio twinning stage, it observes real-world RSRPs $\hat{\mathbf{r}}$ from crowdsourced mobile users and updates the building materials. Specifically, it iteratively calculates the loss function by comparing the real-world RSRPs with the simulated RSRPs from the simulator, and performs multiple gradient descent steps. In the stage of network planning, it iteratively searches the location of the next base station by 1) selecting a base station location (i.e., maximizing the EI acquisition function); 2) querying the DRT to obtain the network performance; and 3) updating the GP surrogate model with new observations.
@@
4Stage 1: Digital radio twin calibrating;
@@ 9   Update parameters $\boldsymbol{\Theta}$ according to (4);
10Keep the checkpoint with the minimum training loss over $E$ epochs, and denote it as $\boldsymbol{\Theta}^{*}$;
@@ V Performance Evaluation
Parameters. We select a $1210~\text{m}\times 1138~\text{m}$ campus area at the University of Nebraska–Lincoln, i.e., the deployment region $\mathcal{A}$. There are $K=127$ objects (i.e., buildings) in this region. We obtain the scene geometry from OpenStreetMap3D and derive the feasible deployment region $\mathcal{F}\subseteq\mathcal{A}$ by extracting the union of building footprints. One existing private 5G base station is located near the southwest corner of $\mathcal{A}$, yielding the initial BS set $\hat{\mathcal{B}}={(\hat{x}_{m},\hat{y}_{m})}_{m=1}^{M}$ with $M=1$. In the region $\mathcal{A}$, we collect $P=1494$ RSRP measurements $\hat{\mathbf{r}}=[\hat{r}_{1},\hat{r}_{2},\ldots,\hat{r}_{P}]$ via drive testing, with a 5G dongle (Quectel RM520N-GL). The accurate location of the receiver is obtained by using RTK-GNSS, providing horizontal accuracy better than $2$ cm. The deployed base stations are equipped with a planar antenna array configured as a $1\times 2$ horizontal linear array with half-wavelength element spacing. The array adopts the 3GPP TR 38.901 directional antenna pattern and vertical polarization, representing a typical 5G small cell configuration. In the evaluation, we aim to select $N=5$ additional BS locations from the feasible region, i.e., $\bar{\mathcal{B}}={(\bar{x}_{n},\bar{y}_{n})}_{n=1}^{5}\subseteq\mathcal{F}$. To calculate coverage, we set $r_{th}=-90\mathrm{dBm}$ and $a=0.2\mathrm{m}$. The default maximum transmission power of the base station is $43\,\mathrm{dBm}$. Additionally, we set $\beta=10$ to balance the contributions of coverage and capacity in the target objective. All simulations were conducted on a desktop with an Intel Core i7 CPU, an NVIDIA RTX 3080 GPU, and Ubuntu 24.04 OS, using Sionna RT v1.1.0 [4]. Specifically, we calculate the RSRP and SNR by using RadioMapSolver on a uniform $0.2\mathrm{m}\times 0.2\mathrm{m}$ grid. To learn the building materials, we use the Adam optimizer [13] with learning rate $\eta=0.01$, for $E=300$ epochs.
@@ V-A Digital Radio Twinning
In this subsection, we evaluate the DRT for all the attributes of fidelity, tractability, and synchronicity, under the AutoPlan algorithm. Specifically, we collect the field measurement via a drive test in the City Campus of the University of Nebraska-Lincoln, as shown in Fig. 5. Fig. 5 and 5 show the learned conductivity and permittivity during DRT calibration.
@@
As shown in Fig. 3, the training MSE loss steadily decreases over the 300 gradient descent steps and converges after around 200 steps, with a minimum value of $113.78$. In other words, the sim-to-real discrepancy is reduced by $11.23(8.98\%)$, by calibrating the parameters of building materials.
@@
In addition, Table I reports the average inference time of the DRT for varying numbers of deployed base stations. Notably, it is $0.39\mathrm{s}$ with a single base station and increases to $1.24,\mathrm{s}$ with five, highlighting the computational cost of large-scale evaluations.